# Graph Condensation: A Survey

### v2

@@ 6 Empirical Studies
TABLE V: The generalizability comparison of GC methods. GC-SNTK is out-of-memory on Reddit. “Avg.” indicates the average value. The highest accuracies are highlighted in bold, and runners-up are underlined.
Dataset Method SGC GCN SAGE APPNP Cheby GAT Avg.
Cora ($r$ = 2.60%) GCond 79.3 80.1 78.2 78.5 76.0 66.2 76.4
GCDM 79.0 77.2 54.4 77.5 72.7 60.4 70.2
GC-SNTK 79.4 81.5 79.1 77.4 73.5 71.9 77.1
SFGC 79.1 81.7 81.9 78.8 79.0 80.8 80.2
Citeseer ($r$ = 1.80%) GCond 70.3 70.6 66.2 69.6 68.3 55.4 66.7
GCDM 72.0 69.5 67.1 69.3 68.7 61.7 68.1
GC-SNTK 71.6 68.4 66.5 69.4 67.6 61.9 67.6
SFGC 71.8 72.4 71.7 70.5 71.8 72.1 71.7
Ogbn-arxiv ($r$ = 0.25%) GCond 63.7 63.2 62.6 63.4 54.9 60.0 61.3
GCDM 61.2 59.6 61.1 62.8 55.4 61.2 60.2
GC-SNTK 62.7 65.1 62.9 62.6 55.1 61.8 61.7
SFGC 64.8 66.1 64.8 63.9 60.7 65.7 64.3
Flickr ($r$ = 0.50%) GCond 46.1 47.1 46.2 45.9 42.8 40.1 44.7
GCDM 44.3 46.8 45.8 45.2 41.8 41.9 44.3
GC-SNTK 45.7 46.8 45.9 45.3 41.3 41.4 44.4
SFGC 47.0 47.0 42.5 40.7 45.4 45.3 44.7
Reddit ($r$ = 0.10%) GCond 89.6 89.6 89.1 87.8 75.5 60.2 82.0
GCDM 88.0 89.7 89.3 88.9 74.9 69.3 83.4
SFGC 89.5 90.0 90.3 88.3 82.8 87.1 88.0
6.1 Evaluation Metric
@@
6.2 Experimental Settings
Methods. We compare four representative GC methods with different optimization strategies: (1) gradient matching-based method: GCond [25]; (2) trajectory matching-based method: SFGC [28]; (3) KRR-based method: GC-SNTK [61]; (4) distribution matching-based method: GCDM [63].
@@ 6.3 Experimental Results
Effectiveness. The condensed graphs generated by GC methods are evaluated to train the same 2-layer GCN [94] for node classification and the test accuracies with standard deviation are reported in Table IV. In the table, “Original graph” refers to the GCN performance which is trained on the original graph and we make the following observations. Firstly, GC methods are not sensitive to the compression rate and a larger condensed graph size does not strictly indicate better performance. Even with extremely small compression rates, GC methods can achieve promising performance. When comparing different optimization strategies, SFGC consistently outperforms other methods across all datasets, primarily due to its long-range trajectory matching. While GC-SNTK utilizes KRR to achieve competitive performance, the graph kernel in KRR is memory-intensive and does not scale well with large-scale graphs.
@@
Generalization. We adhere to the settings commonly used in most GC studies and assess the generalization of GC methods across various GNN models for the node classification task. The GNN models evaluated include GCN [94], SGC [16], SAGE [9], APPNP [96], Cheby [97] and GAT [98]. The detailed accuracies are shown in Table V. Across various GNN architectures, most GC methods (i.e., GCond, GCDM, and SFGC) consistently achieve the best performance on GCN and SGC. This superior performance is attributed to the same convolution kernels [25] in the relay models of these GC methods, which share the same spectral characteristics [64] as those in GCN and SGC. Moreover, SFGC demonstrates a significant improvement over other methods, underscoring the effectiveness of trajectory matching in capturing essential information from the original graph and its superior generalization capabilities.
Efficiency. We report the condensation times of various GC methods on different datasets in Table VI. GCond employs gradient matching to align the original and condensed graphs, necessitating additional back-propagation steps for gradient computation relative to model parameters. In contrast, GCDM modifies the optimization objective and adopts distribution matching, thereby eliminating the need for gradient calculations and achieving faster condensation. Although GC-SNTK utilizes KRR to reduce iterative optimization in the inner loop and achieves rapid condensation on smaller datasets, it struggles to scale to larger graphs. SFGC requires training multiple GNNs on the original graph to provide the trajectories and necessitates multiple updates of the relay model to resemble the training trajectories. All these procedures are computationally intensive and result in extended condensation times.
TABLE VI: The condensation time (seconds) comparison of different GC methods. “OOM” denotes out-of-memory. $r$ is set as 2.60%, 1.80%, 0.25%, 0.50% and 0.10% for five datasets, respectively.
Dataset GCond GCDM GC-SNTK SFGC
Cora 498.9 27.3 30.8 2,524.4
Citeseer 479.3 62.0 19.4 3,877.0
Ogbn-arxiv 14,876.0 690.0 9,353.6 86,288.5
Flickr 1,338.3 209.5 562.6 52,513.0
Reddit 21,426.5 921.9 OOM 246,997.1
@@
Fairness. Following FairGNN [91] and FGD [49], we evaluate GC methods on the Pokec-n dataset to assess group fairness issues within a binary sensitive feature setting. As shown in Table VII, GC methods amplify the graph data bias to varying degrees. GCond and GCDM not only achieve the lowest node classification accuracy but also exhibit the most pronounced bias. More advanced optimization strategies can effectively reduce this bias while preserving predictive performance.
@@
TABLE VII: The comparison of accuracy (ACC, %), AUC (%), demographic parity ($\Delta_{DP}$, %) and equal opportunity ($\Delta_{EO}$, %) for various GC methods.
@@
Metric GCond GCDM GC-SNTK SFGC
Original
graph
ACC $\uparrow$ 51.2±2.1 53.3±1.4 54.0±1.0 54.1±1.1 67.8±0.2
AUC $\uparrow$ 52.2±1.9 53.0±1.6 54.4±1.1 55.5±1.6 72.9±0.2
$\Delta_{DP}\downarrow$ 3.01±1.2 3.3±1.4 2.0±1.7 2.1±1.2 0.2±0.1
$\Delta_{EO}\downarrow$ 3.20±1.0 2.7±1.5 2.6±1.6 2.2±1.1 1.4±1.6
Robustness. Following the RobGC [58], we assess the resilience of GC methods to both random noise and adversarial attacks in the graph structure under the inductive setting. For random noise, this involves adding fake edges and removing real edges from the graph at random. Global adversarial attacks on the structure of the original graph are executed using PR-BCD [99]. Each method’s performance is evaluated at varying noise levels, defined by the ratio of changed edges in the original graph [100]. We assume uniform noise levels across real-world systems, meaning both training and test graphs are subjected to identical noise levels. As illustrated in Figure 10, “Original graph” denotes GNNs trained on the noisy training graph and tested on the noisy test graph. For the GC methods, GNNs are trained on the condensed graphs derived from the noisy training graphs. The adversarial attacks significantly impact classification performance more than random noise at equivalent noise levels. Among the GC methods, GCDM exhibits the lowest accuracy across different noise levels, whereas SFGC demonstrates the most robust performance against both random noise and adversarial attacks. However, as noise levels increase, all performances decline, and the gap between “Original graph” and the various GC methods increases, indicating that noise in the graph structure significantly degrades the quality of the condensed graphs produced.
@@
Figure 10: The performance of GC methods under conditions of random noise and adversarial attacks on Ogbn-arxiv. “Original graph” indicates that GNNs are trained directly on the noisy training graph. In contrast, other methods involve training GNNs on the condensed graph. All GNNs are subsequently evaluated on the noisy test graph.
7 Applications and Resources
@@
7.1 Practical Applications
7.1.1 Hyper-Parameter/Neural Architecture Search
@@
7.1.2 Graph Continual Learning
@@ 7.1.6 Heterogeneous Graphs
7.2 Open-Source Libraries
@@
8 Challenges and Future Directions
@@
8.1 Condensation for Diverse Graphs
@@
8.2 Task-Agnostic Graph Condensation
@@
The optimization strategies currently employed in GC are specifically designed for the classification task, heavily relying on class labels. However, the entanglement of the GC optimization process with task-related labels can potentially hinder performance in other downstream tasks. Therefore, there is a noteworthy potential to develop generalized GC methods for diverse downstream tasks, including link prediction [112], anomaly detection [113], graph reconstruction [114], and graph generation [115], etc. Furthermore, considering the extensive time and resources required for data annotation, the development of self-supervised GC can significantly advance and expand the adaptability of GC techniques.
8.3 Efficient Evaluation for Condensed Graphs
@@
In the realm of GC, the primary objective is to create high-quality condensed graphs. However, a systematic and well-established methodology for assessing the condensed graph quality is notably absent in the field [28]. This gap is particularly critical given the complex optimization involved in GC. The assessment of the condensed graph quality is essential not only for monitoring the condensation progress but also for potentially streamlining the optimization through early-stop strategies. Currently, GC practices leverage the downstream GNNs’ performance as the indicator for assessment. However, this method adds considerable computational overhead to the GC process. Moreover, the concept of quality in condensed graphs is inherently complex, varying significantly across different applications and necessitating a multifaceted approach to evaluation. Considering these challenges, there is a pressing requirement to develop comprehensive and efficient evaluation metrics. Such metrics would not only accelerate the condensation process but also broaden the range of GC applications.
8.4 Secure Graph Condensation
@@
The condensation of graphs not only facilitates efficient training of GNNs but also contributes significantly to reducing the storage overhead and accelerating data transmission. However, the security of highly compressed graph datasets to various privacy threats [116] remains under-explored. Recent study [88] has exposed vulnerabilities of condensed graphs to backdoor attacks, while additional threats such as adversarial attacks and inference breaches present substantial challenges for deploying GC in security-sensitive sectors, including confidential communications, financial systems, and industrial controls. In light of these challenges, the development of secure, privacy-preserving GC algorithms becomes imperative.
8.5 Explainable Graph Condensation
@@
Another promising avenue for GC research is the development of explainable GC methods. Current GC techniques, while efficient in reducing graph size, often operate as a black box, offering limited insight into how condensed graphs represent the original graph data [117]. An explainable condensation can illuminate the intricate patterns within the graph, thereby enhancing user trust and acceptance. This is particularly important in sectors like healthcare [118] and finance [33], where the interpretability and justification of decision-making processes are critical. Consequently, by elucidating the logic behind data transformations, explainable GC has the potential to open up new avenues for its application, establishing GC as a more reliable and transparent tool in data-centric technology fields.
9 Conclusion

### v3

@@ VI Empirical Studies
VI-A Evaluation Metric
@@
TABLE V: The generalizability comparison of GC methods on Ogbn-arxiv dataset. “Opt.” and “Avg.” indicate the optimization strategy and average value, respectively. The highest accuracies are highlighted in bold, and runners-up are underlined. $r$ is set as 0.25%.
Opt.
Method SGC GCN SAGE APPNP Cheby GAT Avg.
GM GCond 63.7 63.2 62.6 63.4 54.9 60.0 61.3
GCond-X 64.7 64.2 64.4 61.5 59.5 60.1 62.4
DosCond 63.3 63.5 62.1 63.5 55.1 60.4 61.3
SGDD 64.3 65.8 63.4 63.3 55.9 61.4 62.4
DM GCDM 61.2 59.6 61.1 62.8 55.4 61.2 60.2
GCDM-X 64.4 61.2 63.4 60.5 60.2 60.0 61.6
SimGC 64.3 66.4 60.4 61.5 54.7 61.1 61.4
KRR GC-SNTK 62.7 65.1 62.9 62.6 55.1 61.8 61.7
GC-SNTK-X 64.0 65.5 62.4 61.8 58.7 60.9 62.2
TM SFGC 64.8 66.1 64.8 63.9 60.7 65.7 64.3
GCSR 65.6 65.4 65.4 64.4 58.9 63.5 63.9
VI-B Experimental Settings
Methods. To assess the efficacy of various optimization strategies, we evaluate 11 representative GC methods categorized under 4 distinct optimization strategies: (1) gradient matching (GM)-based method: GCond, GCond-X [26], DosCond [54] and SGDD [31]; (2) distribution matching (DM)-based method: GCDM, GCDM-X [65] and SimGC [70]; (3) KRR-based method: GC-SNTK and GC-SNTK-X [63]; (4) trajectory matching (TM)-based method: SFGC [29] and GCSR [62]. Notice that the suffix “-X” represents the graphless variant.
@@ VI-C Experimental Results
Effectiveness. The condensed graphs generated by GC methods are evaluated to train the same 2-layer GCN [97] for node classification and the test accuracies with standard deviation are reported in Table IV. In the table, “Original graph” refers to the GCN performance which is trained on the original graph and we make the following observations. (1) GC methods are not sensitive to the compression rate and a larger condensed graph size does not strictly indicate better performance. Even with extremely small compression rates, GC methods can achieve promising performance. (2) The graphless variants demonstrate comparable performance to the method with the structure modeling, indicating the graph structure information can be effectively encoded into the condensed nodes. (3) When comparing different optimization strategies, trajectory matching-based methods consistently show superior performance, outperforming other strategies across various datasets. While KRR-based methods achieve competitive performance, the graph kernel in KRR is memory-intensive and does not scale well with large-scale graphs. Distribution matching-based methods perform well on smaller datasets, whereas gradient matching-based methods are more effective on larger datasets.
Figure 10: The condensation time (seconds) comparison of different GC methods on Ogbn-arxiv dataset. $r$ is set as 0.25%.
@@
Generalization. We adhere to the settings commonly used in most GC studies and assess the generalization of GC methods across various GNN models for the node classification task. The GNN models evaluated include GCN [97], SGC [17], SAGE [10], APPNP [99], Cheby [100] and GAT [101]. The detailed accuracies are shown in Table V. Across various GNN architectures, most GC methods consistently achieve the best performance on GCN and SGC. This superior performance is attributed to the same convolution kernels [26] in the relay models of these GC methods, which share the same spectral characteristics [66] as those in GCN and SGC. Among different optimization strategies, trajectory matching-based methods demonstrate a significant improvement over other methods, underscoring the effectiveness of trajectory matching in capturing essential information from the original graph and its superior generalization capabilities.
Efficiency. We report the condensation times of various GC methods in Fig. 10 and make the following observations. (1) Distribution matching achieves the most efficient condensation process among the optimization strategies, primarily due to its elimination of the need for gradient calculations. In contrast, trajectory matching and gradient matching are computational intensive. (2) Without graph structure modeling, graphless variant methods expedite the condensation process. (3) The one-step matching strategy used in DosCond eliminates the intensive model updates required in the inner loop while maintaining competitive classification performance.
@@
Fairness. Following FairGNN [94] and FGD [51], we evaluate the representative GC method in each optimization strategy on the Pokec-n dataset, and assess group fairness issues within a binary sensitive feature setting. As shown in Table VI, GC methods amplify the graph data bias to varying degrees. Gradient matching and distribution matching not only achieve the lowest node classification accuracy but also exhibit the most pronounced bias. More advanced optimization strategies, i.e., KRR and trajectory matching, can effectively reduce this bias while preserving predictive performance.
@@
TABLE VI: The comparison of AUC (%), demographic parity ($\Delta_{DP}$, %) and equal opportunity ($\Delta_{EO}$, %) for GC methods with various optimization strategies.
Opt. GM DM KRR TM Origianl graph
@@
Method GCond SGDD GCDM SimGC GC-SNTK SFGC GCSR
AUC $\uparrow$ 52.2±1.9 55.1±0.9 53.0±1.6 55.3±1.3 54.4±1.1 55.5±1.6 54.9±0.6 72.9±0.2
$\Delta_{DP}\downarrow$ 3.0±1.2 2.5±1.0 2.3±1.4 2.4±1.1 2.0±1.7 2.1±1.2 1.9±0.9 0.2±0.1
$\Delta_{EO}\downarrow$ 3.2±1.0 3.0±1.1 2.7±1.5 2.8±1.2 2.6±1.6 2.2±1.1 2.0±1.2 1.4±1.6
Robustness. We evaluate the robustness of GC methods to structural noise, which involves randomly adding and removing edges. The noise level is set at 100%, defined by the proportion of modified edges in the original graph [102]. As illustrated in Table VII, the noisy graph structure significantly impacts the classification performance of models trained on condensed graphs. The performance disparity between noisy condensed graphs and noisy original graphs is greater than that observed with clean graphs, suggesting that structural noise severely affects the quality of condensed graphs. Among the GC methods, distribution matching is most sensitive to random noises, while trajectory matching demonstrates robust performances.
@@
TABLE VII: The performance of GC methods against random structure noise on Ogbn-arxiv. Noise level is set as 100%. “Original graph” indicates that GNNs are trained on the noisy original graph. In contrast, other methods involve training GNNs on the condensed graph.
Method GM DM KRR TM Origianl graph
GCond SGDD GCDM SimGC GC-SNTK SFGC GCSR
Clean graph 63.2±0.3 65.8±1.2 59.6±0.4 66.4±0.3 65.1±0.8 66.1±0.4 65.4±0.8 71.4±0.1
Noisy graph 50.1±0.5 52.9±0.9 49.4±0.7 51.7±0.4 52.7±0.7 54.6±0.7 55.5±0.5 61.1±0.1
VII Applications and Resources
@@
VII-A Practical Applications
VII-A1 Hyper-Parameter/Neural Architecture Search
@@
VII-A2 Graph Continual Learning
@@ VII-A6 Heterogeneous Graphs
VII-B Open-Source Libraries
@@
VIII Challenges and Future Directions
@@
VIII-A Condensation for Diverse Graphs
@@
VIII-B Task-Agnostic Graph Condensation
@@
The optimization strategies currently employed in GC are specifically designed for the classification task, heavily relying on class labels. However, the entanglement of the GC optimization process with task-related labels can potentially hinder performance in other downstream tasks. The recent study, CTGC [113], investigates self-supervised learning to enhance the task adaptability of GC for link prediction, and clustering tasks. This confirms the feasibility of extending GC to a wider spectrum of practical scenarios and delineates promising directions for future research, including anomaly detection [114], graph reconstruction [115], and graph generation [116].
VIII-C Efficient Evaluation for Condensed Graphs
@@
In the realm of GC, the primary objective is to create high-quality condensed graphs. However, a systematic and well-established methodology for assessing the condensed graph quality is notably absent in the field [29]. This gap is particularly critical given the complex optimization involved in GC. The assessment of the condensed graph quality is essential not only for monitoring the condensation progress but also for potentially streamlining the optimization through early-stop strategies. Currently, GC practices leverage the downstream models, i.e., GNN [26] or graph neural tangent kernel [29], as the indicator for assessment. However, this method adds considerable computational overhead to the GC process. Moreover, the concept of quality in condensed graphs is inherently complex, varying significantly across different applications and necessitating a multifaceted approach to evaluation. Considering these challenges, there is a pressing requirement to develop comprehensive and efficient evaluation metrics. Such metrics would not only accelerate the condensation process but also broaden the range of GC applications.
VIII-D Secure Graph Condensation
@@
The condensation of graphs not only facilitates efficient training of GNNs but also contributes significantly to reducing the storage overhead and accelerating data transmission. However, the security of highly compressed graph datasets to various privacy threats [117] remains under-explored. The recent study [91] has exposed vulnerabilities of GC and injected malicious information into condensed graphs to backdoor the GNNs trained on them. This highlights concerns over additional threats, such as adversarial attacks [60] and inference breaches [118], which pose significant challenges for deploying GC in security-sensitive domains, including confidential communications, financial systems, and industrial controls. In light of these challenges, the development of secure, privacy-preserving GC algorithms becomes imperative.
VIII-E Explainable Graph Condensation
@@
Another promising avenue for GC research is the development of explainable GC methods. A relevant work [56] utilizes explainable algorithms to assess the importance of condensed nodes and prune the condensation redundancy. However, this method still provides limited insight into how condensed graphs represent the original data [119]. An explainable condensation can illuminate the intricate patterns within the graph, thereby enhancing user trust and acceptance. This is particularly important in sectors like healthcare [120] and finance [36], where the interpretability and justification of decision-making processes are critical. Consequently, by elucidating the logic behind data transformations, explainable GC has the potential to open up new avenues for its application, establishing GC as a more reliable and transparent tool in data-centric technology fields.
IX Conclusion